OpenStudy (anonymous):

can someone help me with 7th grade algebra? like, give me some problems?

OpenStudy (anonymous):

look for any algrbraic text itll help you!!!!

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Here are a few I recall from that age: Well my first (and easiest) question was to find all integer triples (x,y,z), and where n is an integer > 2, such that \[ x^n + y^n = z^n \] A slightly tougher one was: Prove that the set A defined as \[ \{ z | z \in \mathbb{C}, \zeta (z) = 0, \zeta (s) := \sum 1/n^s \} \] is a subset of the set B-union-C, where B defined as \[ \{ z | z = 0.5 + xi, x \in \mathbb{R} \} \] and C is defined as \[ \{ -2n | n \in \mathbb{N} \} \]

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

i have no idea what any of that means

OpenStudy (anonymous):

:/